OpenStudy (anonymous):

write the equation of the circle in standard form. x^2 + y^2 + 12x +10y +61 = 16

OpenStudy (anonymous):

Complete the squares.\[x^2 + y^2 + 12x +10y +61 = 16 \implies(x^2+ 12x+36) + (y^2 +10y +25)= 16\]\[\implies (x+6)^2+(y+5)^2=16\]

OpenStudy (anonymous):

Looks like the center is at (h,k)=(-6,-5) with r=4.

OpenStudy (anonymous):

Thank you very much.